# Nonequilibrium physics of brain dynamics

### v1

@@ 1.2.1 The equilibrium Ising model
In a pair of seminal papers, E.T. Jaynes introduced the principle of maximum entropy [55, 56]. In simple terms, the principle argues that given empirical data and a set of constraints to which the model/distribution which generated the data much adhere to, the distribution which satisfies the constraints yet maximises the entropy/uncertainty is the preferred choice [57]. A classic example is to consider the binary activity of a set of neurons, $(s_{1},...,s_{N})(t_{k})$, each being active $(+1)$ or inactive $(-1)$ at each time-step $t_{k}$. From observed data in the form of neural spike-trains, one can calculate both the average firing probability of each neuron, $\langle s_{i}\rangle$, as well as the pairwise correlations, $\langle s_{i}s_{j}\rangle-\langle s_{i}\rangle\langle s_{j}\rangle$. Under these constraints, the distribution that maximises the entropy is precisely the Boltzmann distribution,
@@ 1.2.2 Symmetric whole-brain models
where $x_{i}(t)$ represents the activity of region $i$ over time, $f_{i}(\cdot)$ represents a nodal dynamic, $\bm{A}=(A_{ij})$ represents the pairwise regional couplings and $g_{ij}(\cdot)$ represents a coupling function [74, 75]. Typically, regions are assumed to be homogeneous ($f=f_{i},g=g_{ij}$), with regional dynamics given by neural-mass or oscillatory models [5], such as the Hopf [76] or dynamic mean field models [77]. The interaction network $\bm{A}$ is the structural connectivity matrix (SC) between brain regions, usually a participant-averaged network obtained separately from the same functional imaging that generated the time-series [75]. Although in animals invasive techniques such as tract-tracing can measure directed, asymmetric connections between brain regions [78], no such data is directly available for human brain networks, so-called connectomes, which rely on non-invasive imaging techniques [79], such as diffusion tensor imaging (DTI), which lead to undirected representations of brain networks ($A_{ij}=A_{ji}$, $\bm{A}=\bm{A}^{\top}$) despite the fact that they are inherently directed [80].
@@
where $r(x,y)$ represents the Pearson correlation coefficient (PCC) between the series $x$ and $y$ [84]. As a result, the FC is typically symmetric ($C_{ij}=C_{ji}$, $\bm{C}=\bm{C}^{\top}$) and therefore incapable of capturing asymmetries in interaction strengths. Between the symmetric models and the symmetric summary statistics, connectome-based whole-brain models typically fail to capture the nonequilibrium dynamics of the underlying data. As a result, one must turn to simplified or statistical models such as those described in Sec. 3.111It is worth noting that whilst PCC is the standard metric, FC can be defined using any pairwise metric of which there are many, including a range of directed metrics [86]. It is unclear if symmetric models can fit asymmetric summary statistics.
@@ 1.2.3 Directedness in non-human connectomes
Whilst SC in the human brain is typically restricted to undirected representations, invasive methods have reconstructed directed connectomes for animal and insect brains [78, 80]. For example, the FlyWire project recently published a complete, directed wiring-diagram of the Drosophila brain [87]. Its directed structure reveals that some neurons operate as ‘broadcasters’, with many more out-going connections than incoming, whilst others are ‘integrators’ with a larger number of incoming connections [88]. Moreover, the density of reciprocal and non-reciprocal connections vary significantly throughout the fly brain.
In addition to the fly, directed connectomes are available for the macaque [78], cat [89], mouse [90] and C. Elegans [91]. Directionality has significant impact on the structural properties of these networks such as degree-distribution, clustering coefficient and small-world index [80], with rich-club regions showing higher in-degree than out-degree [80, 92]. Finally, directed structure can constrain dynamics with a strong correlation being shown between in-degree characteristics and resting-state fMRI in the mouse [93].
@@ 2.1 Continuous-space processes and Langevin equations
[…]
@@
[…]
@@
[…]
@@ 2.2.1 Continuous-time processes
A stochastic process defined on a finite discrete state-space describes the (time-)evolution of a sequence of random variables which take values in a discrete support, the set of possible states. Such a process is defined by its probability distribution function, $P({X}={x},t)$, where $t$ is time, ${X}$ is the state of the process and ${x}$ is a value in the support. Without loss of generality, we consider the support to be a finite set of natural numbers, $\{1,...,n\}$. Therefore, we can consider $\bm{P}(t)$ to be the $n$-dimensional process with $P_{i}(t)$ being the probability of being in state $i$ at time $t$. Typically, one assumes that the process is Markovian meaning that the dynamics only depend on the current state of the system and are memoryless. In this case, $\bm{P}(t)$ can be described by the master equation,
@@
where $\bm{Q}(t)$ is the generator matrix [127]. Another typical assumption is that the generator matrix is constant in time, i.e. $\bm{Q}(t)\equiv\bm{Q}$, and defines a continuous-time Markov chain (CTMC). The off diagonal entries $q_{ab}$ of $\bm{Q}$ represent the transition intensity from state $b$ to $a$ whilst diagonal entries are defined to be $q_{aa}=-\sum_{b}q_{ba}$, thus each column sums to 0 which ensures that probability is conserved in time [127]. Master equation dynamics assume that transitions occur after random, state-dependent waiting-times which are exponentially-distributed with rate parameter $-q_{aa}$ in state $a$.
@@ 3 Analysis of neuroimaging and continuous-valued recordings
Large-scale neural activity can be recorded using invasive and non-invasive neuroimaging techniques such as functional magnetic resonance imaging (fMRI), magnetoencephalography (MEG) and electroencephalography (EEG). Such data is often processed into a MVTS where each real-valued variable represents the activity of a discrete brain region over time, subject to a specific parcellation of the brain into distinct regions [75]. The dimension of neuroimaging data can vary drastically depending on modality and experimental paradigm, with fMRI being particularly spatially precise, but temporally imprecise, and MEG being particularly spatially imprecise with greater temporal precision. Typical parcellations of the human brain can vary between $N\approx 10^{0}-10^{6}$. In high-dimensions, many traditional techniques for the analysis of stochastic dynamics such as force field reconstruction [134, 135] cannot be applied. Instead, one must opt for either model-based approaches, where one assumes the functional form of the dynamical system, or model-free measures which are typically correlates of the overall level of nonequilibrium.
@@ 3.1.2 The Kuramoto model
[…]
@@ 3.1.3 The asymmetric Hopf model
for $j=1,...,N$, where $z_{j}=x_{j}+\bm{i}y_{j}$ is the complex-valued signal for node $j$, $\omega_{j}$ is the intrinsic regional frequency, $a_{j}$ is a bifurcation parameter, $\bm{A}$ is the network of interactions and $\eta_{j}$ is additive noise that is uncorrelated in time and between nodes. In the absence of noise and interactions, for $a_{j}<0$, the dynamics converge to a stable fixed point. The system goes through a bifurcation at $a_{j}=0$ after which the dynamics then converge to the stable limit-cycle. The model can also be written in Cartesian coordinates,
@@
for $j=1,...,N$, where, in both cases, $x_{j}(t)$ is considered to be the activity of region $j$ over time.

### v2

@@ 1 Introduction
1.1 Neurons, populations and neural dynamics
It has long been known that the brain is the seat of cognition for both humans and many other animals, as such it was of profound interest to early philosophers and scientists alike. It was not until the work of Golgi and y Cajal, that neurons were identified as the principals computational unit of the brain [33, 34]. Despite the zoo of different cell types in the brain, the study of cognition has since focused primarily on neurons and their connectivity.
Briefly, neurons communicate via electrical signals that propagate along their axons (Panel A, Fig. 2). They operate as threshold units, requiring a sufficient input of current to prompt a burst of activity known as an action potential or spike that travels along the axon. Hodgkin and Huxley were able to derive the equations for neuronal spiking from the dynamics of ionic channels in the neuron [4], thus allowing for the modelling of neuronal activity. These models take the form of ordinary differential equations (ODEs) and have a range of dynamic regimes with different spiking behaviour (Panel B, Fig. 2). Through synapses, neurons are connected and action potentials from incoming connections serve as input current for a post-synaptic neuron, thus potentially inducing further spikes. In this fashion, populations of neurons exhibit coordinated, collective behaviour. The activity of a neuronal network can be simplified by discretizing the spikes as discrete threshold events, ignoring the continuous dynamics of the action potential (Panel C, Fig. 2). The dominant connectionist paradigm in neuroscience argues that cognition in the brain is performed in a distributed fashion amongst networks of simple processing units, and has inspired the development of artificial neural networks and deep learning [35].
Figure 2: Neurons, populations and dynamics. A. A schematic diagram of a single neuron. Adapted from Notjim by $C$ BY-SA 3.0 (https://commons.wikimedia.org/w/index.php?curid=4824168). B. Action potentials from different dynamic regimes. These are simulated using the Izhikevich model [2]. C. Neurons are organised into networks of connected units, which produce collective activity. This can be visualised as a spike-train in a raster plot. Drawing by Santiago Ramón y Cajal (Public Domain). D. The average activity of a population or network or neurons can be approximated with a mean-field model. These models have an array of dynamic regimes. These are simulated from the Jansen-Rit neural-mass model [5].
Recording from individual neurons in the brain is challenging and requires invasive electrophysiology. As such, these modalities primarily focus on small animals such as mice and rats, with limited data for humans. On the other hand, neuroimaging and encephalography methods, such as functional magnetic resonance imaging (fMRI), magnetoencephalography (MEG) and electroencephalography (EEG), are able to non-invasively capture large-scale activity in the brain. These modalities record proxies of neural activity such as blood-flow into brain regions (fMRI), magnetic fields measurable from outside the head (MEG), and electrical potentials on the scalp (EEG). As a result, these data can be considered to be the average activity in a region of interest (ROI) of the brain, known as a neural-mass. This neural-mass is composed of a large number of interacting neurons with heterogenous connectivity and cell-types. Nevertheless, for the purposes of modelling such data, we typically opt for a mean-field approach, where a single model is used to describe the average activity of a entire neural population. A diverse range of neural-mass models exist, with a variety of complexity, and they are able to reproduce an array of different activity types, in alignment with neuroimaging data [5] (Panel D, Fig. 2). By considering macroscopic connections between neural masses, white-matter fibres visible under diffusion tensor imaging (DTI) [36], we can model the collective behaviour of the whole brain, constrained by this connectivity, up to the resolution of the data at hand.
Both the connectionist perspective on neuronal activity, and the resolution of neuroimaging, encephalography and electrophysiological data, dictate the level of abstraction from which we attempt to understand neural systems. Modelling the brain as an interconnected network of discrete interacting units is the dominant paradigm that is most aligned with experimental data - although this is performed at a diversity of scales, including both discrete and continuous outputs. This naturally sets the scene for a mathematical paradigm that focuses on nonlinear dynamics, network science, statistical physics, and time-series analysis.
@@ 1.2 Nonequilibrium signatures of consciousness and cognition
1.3 Comments on consciousness and its correlates
Before moving on from the previous section, it is worth clarifying our use of the often-controversial term consciousness. In simple terms, being conscious refers to the subjective phenomenon of ‘having an experience’. Despite broad agreement that we are referring to the same concept, consciousness remains an illusory and vague concept that so far has eluded a satisfactory rigorous scientific definition. Nevertheless, developing a mathematical or scientific theory of consciousness is often described as one of the most important, interesting and challenging problems of science and philosophy, especially by the people who write extensively about it.
Despite many valiant efforts, the study of consciousness is, as of yet, far from being an empirical science due to the challenges of developing consensus definitions or relating theory to empirical analysis. Moreover, the paradox of the ‘hard problem of consciousness’, as posited by Chalmers [59], suggests that even a complete neuroscientific understanding of consciousness may fail to give a satisfactory answer to the more esoteric question of why we have experiences at all.
Despite a clear media appeal, such discussions are beyond the scope of our current mathematical and physical frameworks, and we make no claims about the relationship between nonequilibrium physics and the hard problem. Nevertheless, neuroimaging collected from participants in a diverse array of brain states can give way to markers that indicate the level of activity or awareness that are usually conflated with the notion of consciousness. For example, most would agree that a cadaver, an anaesthetised person, a sleeping person, and an awake person can be placed on a scale of increasing levels of consciousness, even if this appears impossible to objectively confirm. Proceeding under this assumption, we can probe the neural dynamics underlying these states to investigate dynamical features that display a correlation with the assumed ladder of conscious states. These so-called ‘neural correlates of consciousness’ [60], which we refer to in the proceeding section, make no direct headway towards the hard problem of consciousness, yet may give us insight into the dynamical features that appear to orchestrate conscious behaviour.
It is from this vantage point that we make reference to the EPR or level of nonequilbrium as a signature of conscious behaviour [46], arguing that any more direct claims about the application of nonequilibrium physics to an elusive ‘theory of consciousness’ goes beyond the evidence collected hereto.
@@ 1.4.1 The equilibrium Ising model
In a pair of seminal papers, E.T. Jaynes introduced the principle of maximum entropy [61, 62]. In simple terms, the principle argues that given empirical data and a set of constraints to which the model/distribution which generated the data much adhere to, the distribution which satisfies the constraints yet maximises the entropy/uncertainty is the preferred choice [63]. A classic example is to consider the binary activity of a set of neurons, $(s_{1},...,s_{N})(t_{k})$, each being active $(+1)$ or inactive (1)(-1) at each time-step $t_{k}$. From observed data in the form of neural spike-trains, one can calculate both the average firing probability of each neuron, $\langle s_{i}\rangle$, as well as the pairwise correlations, $\langle s_{i}s_{j}\rangle-\langle s_{i}\rangle\langle s_{j}\rangle$. Under these constraints, the distribution that maximises the entropy is precisely the Boltzmann distribution,
@@ 1.4.2 Symmetric whole-brain models
where $x_{i}(t)$ represents the activity of region $i$ over time, $f_{i}(\cdot)$ represents a nodal dynamic, $\bm{A}=(A_{ij})$ represents the pairwise regional couplings and $g_{ij}(\cdot)$ represents a coupling function [80, 81]. Typically, regions are assumed to be homogeneous ($f=f_{i},g=g_{ij}$), with regional dynamics given by neural-mass or oscillatory models [5], such as the Hopf [82] or dynamic mean field models [83]. The interaction network $\bm{A}$ is the structural connectivity matrix (SC) between brain regions, usually a participant-averaged network obtained separately from the same functional imaging that generated the time-series [81]. Although in animals invasive techniques such as tract-tracing can measure directed, asymmetric connections between brain regions [84], no such data is directly available for human brain networks, so-called connectomes, which rely on non-invasive imaging techniques [36], which lead to undirected representations of brain networks ($_{ij}=A_{ji}, 𝑨=𝑨\bm$$\bm{A}^{\top}$) despite the fact that they are inherently directed [85].
@@
where $r(x,y)$ represents the Pearson correlation coefficient (PCC) between the series $x$ and $y$ [89]. As a result, the FC is typically symmetric ($_{ij}=C_{ji}, 𝑪=𝑪\bm$$\bm{C}^{\top}$) and therefore incapable of capturing asymmetries in interaction strengths. Between the symmetric models and the symmetric summary statistics, connectome-based whole-brain models typically fail to capture the nonequilibrium dynamics of the underlying data. As a result, one must turn to simplified or statistical models such as those described in Sec. 3.111It is worth noting that whilst PCC is the standard metric, FC can be defined using any pairwise metric of which there are many, including a range of directed metrics [91]. It is unclear if symmetric models can fit asymmetric summary statistics.
@@ 1.4.3 Directedness in non-human connectomes
Whilst SC in the human brain is typically restricted to undirected representations, invasive methods have reconstructed directed connectomes for animal and insect brains [84, 85]. For example, the FlyWire project recently published a complete, directed wiring-diagram of the Drosophila brain [92]. Its directed structure reveals that some neurons operate as ‘broadcasters’, with many more out-going connections than incoming, whilst others are ‘integrators’ with a larger number of incoming connections [93]. Moreover, the density of reciprocal and non-reciprocal connections vary significantly throughout the fly brain. Simulated dynamics on the asymmetric fly brain shows violation of the fluctuation-dissipation theorem, and the emergence of nonequilibrium dynamics [94].
In addition to the fly, directed connectomes are available for the macaque [84], cat [95], mouse [96] and C. Elegans [97]. Directionality has significant impact on the structural properties of these networks such as degree-distribution, clustering coefficient, small-world index and scale-free properties [85, 98], with rich-club regions showing higher in-degree than out-degree [85, 99]. Finally, directed structure can constrain dynamics with a strong correlation being shown between in-degree characteristics and resting-state fMRI in the mouse [100].
@@
1.5 Relating model properties with empirical data
Throughout this review, we consider abstract models, typically originating from the modelling of physical systems, and apply them to neuroscientific data. It is worth taking a moment to consider the relationship between the properties of a mathematical model and the data which it attempts to describe. To begin with, we note that within the paradigm of statistical physics, such a problem runs rampant. This problem is a natural consequence of the development of statistical physics to describe physical systems such as gases, magnets or fluids, only to find that the same principles apply to a range of systems, both abstract and (bio)physical. A classical example is the relationship between Gibbs’ notion of entropy for a collection of classical particles [113] and Shannon’s information-theoretic entropy [114]. The first is a thermodynamic, and therefore physical, property of the system which describes the number of microstates consistent with a collection of macroscopic thermodynamic properties. In Shannon’s form, it can readily be applied to abstract systems to describe the uncertainty in a random variable given its probability distribution. Shannon’s entropy is precisely the same as Gibbs’ when the random variable describes the energy of the microstates - which need not be the case.
In a similar vein, throughout statistical physics approaches to neuroscience or other complex systems, one may make reference to the temperature, energy, and entropy production rate of a system, without necessarily referring to the physical properties of the system, but instead describing the probabilistic behaviour of its dynamics at a particular coarse-grained scale [115, 116]. This distinction is most important when describing a system at an abstract level, but which is physically instantiated in the world. For example, whilst we may describe the ‘temperature of activity in the brain’ through an analogy with the Ising model, this may or may not be related to the physical temperature of the brain in real terms - the relationship is entirely unclear. Throughout this review, we focus on the EPR of neural systems, in reference to the irreversibility of neural dynamics and the presence of dissipative forces in phase-space. The relationship between the EPR of a neural model or data and its physical properties is not apparent. The thermodynamic properties of information processing [32, 117] complicate this analogy even further, by introducing another notion of entropy production for computations - including those being performed in the brain. With the exception of Sec. 5, we refer to the abstract EPR associated with the dissipative phase-space dynamics, and venture no claims about the energetic costs of either the brain as an organ, or the brain as a computer.
A closely connected problem is determining when data can be said to take on the property of a dynamic model. For example, when a model that exhibits criticality, chaos, or turbulence - each a well-defined concept in mathematical and physical systems - is fit to neural data, where these phenomena are less well-defined, is it fair to say that the brain is then critical, chaotic, or turbulent? It remains a point of contention. When addressing a similar topic, Bassett et al [118] introduce three notions of validity for a network-based model of the brain, which we adapt here to notions of neural dynamics.
  1. 1.
    Descriptive validity: Does the model produce trajectories that resemble the statistical properties of the data?
  2. 2.
    Explanatory validity: Does the model provide evidence for a causal mechanism in the system? By extension, can it be used to test causal relationships on the basis of this model?
  3. 3.
    Predictive validity: Can the model be used to make predictions that are later confirmed by empirical data?
It is clear from these definitions that the lowest barrier to entry for a neural model is descriptive validity which simply describes whether or not a model can be tuned to fit neural data with sufficient accuracy. In this case, attributing properties of the model to the data is the least relevant and supported. In our context, this could be related to the finding that neural dynamics exhibit significant [^2] See Sec. 3.5 for a discussion of significance testing for time-irreversibility. irreversibility, and are therefore best described by nonequilibrium models. On the other hand, linking nonequilibrium dynamics to the hierarchical asymmetric structure of effective brain networks then provides a causal mechanism for the violation of detailed balance [30]. Similarly, the leveraging of nonequilibrium dynamics for sequencing in associative memory models [79] is also a causal, explanatory mechanism. Finally, showing that the level of nonequilibrium varies between brain states allows for predictive validity, where nonequilibrium signatures can be used to classify unlabelled samples from neuroimaging data at a variety of consciousness levels.
In the case of criticality, chaos, or turbulence, a similar analysis can be made. If the optimal working point of the model coincides with the specific regime associated with a property, this implies descriptive validity. The explanatory validity depends on the causal mechanism associated with the property i.e. optimal sensitivity to information for critical dynamics. Ideally this hypothesis is made apriori, rather than an aposteriori explanation of observed results. Finally, if the causal mechanism allows for predictions to be made about the dynamics in unseen data, or the behaviour of the system in response to perturbations, it satisfies predictive validity.
@@ 2.1 Continuous-space processes and Langevin equations
where $\nabla\cdot$ represents the divergence in space and $\nabla$ is the gradient. Over time, the system can converge to a steady-state specified by a stationary density, $P_{\text{ss}}(\mathbf{x})$, which no longer depends on time. The stationary probability flux therefore has zero divergence, $\nabla\cdot\bm{J}_{ss}\equiv 0$. In an equilibrium steady-state (ESS) the probability flux vanishes, $\bm{J}_{ss}\equiv 0$, which is the detailed balance condition for SDEs. On the other hand, if $\bm{J}_{\text{ss}}\neq 0$, the steady-state contains probability currents, violates detailed balance, and is called a nonequilibrium steady-state (NESS).
@@
$\dot{S}(t)$ $=\int_{\mathbb{R}^{n}}(\log P(\mathbf{x},t)+1)\nabla\cdot\bm{J}(\mathbf{x},t)\;d\mathbf{x},$ (11)
@@
$\dot{S}(t)$ $=-\int_{\mathbb{R}^{n}}\nabla(\log P(\mathbf{x},t)+1)\cdot\bm{J}(\mathbf{x},t)\;d\mathbf{x},$ (12)
$=-\int_{\mathbb{R}^{n}}\frac{\nabla P(\mathbf{x},t)}{P(\mathbf{x},t)}\cdot\bm{J}(\mathbf{x},t)\;d\mathbf{x}.$ (13)
@@ 2.2.1 Continuous-time processes
A stochastic process defined on a finite discrete state-space describes the (time-)evolution of a sequence of random variables which take values in a discrete support, the set of possible states. Such a process is defined by its probability distribution function, P(X=x,t)P({X}={x},t), where $t$ is time, X{X} is the state of the process and x{x} is a value in the support. Without loss of generality, we consider the support to be a finite set of natural numbers, $\{1,...,n\}$. Therefore, we can consider $\bm{P}(t)$ to be the $n$-dimensional process with $P_{i}(t)$ being the probability of being in state $i$ at time $t$. Typically, one assumes that the process is Markovian meaning that the dynamics only depend on the current state of the system and are memoryless. In this case, $\bm{P}(t)$ can be described by the master equation,
@@
where $\bm{Q}(t)$ is the generator matrix [140]. Another typical assumption is that the generator matrix is constant in time, i.e. $\bm{Q}(t)\equiv\bm{Q}$, and defines a continuous-time Markov chain (CTMC). The off diagonal entries $q_{ab}$ of $\bm{Q}$ represent the transition intensity from state $b$ to $a$ whilst diagonal entries are defined to be $q_{aa}=-\sum_{b}q_{ba}$, thus each column sums to 0 which ensures that probability is conserved in time [140]. Master equation dynamics assume that transitions occur after random, state-dependent waiting-times which are exponentially-distributed with rate parameter $q_{aa}$ in state $a$.
@@ 3 Analysis of neuroimaging and continuous-valued recordings
Large-scale neural activity can be recorded using invasive and non-invasive neuroimaging techniques such as fMRI, MEG and EEG. Such data is often processed into a MVTS where each real-valued variable represents the activity of a discrete brain region over time, subject to a specific parcellation of the brain into distinct regions [81]. The dimension of neuroimaging data can vary drastically depending on modality and experimental paradigm, with fMRI being particularly spatially precise, but temporally imprecise, and MEG being particularly spatially imprecise with greater temporal precision. Typical parcellations of the human brain can vary between $N\approx 10^{0}-10^{6}$. In high-dimensions, many traditional techniques for the analysis of stochastic dynamics such as force field reconstruction [147, 148] cannot be applied. Instead, one must opt for either model-based approaches, where one assumes the functional form of the dynamical system, or model-free measures which are typically correlates of the overall level of nonequilibrium.
@@ 3.1.2 The Kuramoto model
$Re^{\bm{i}\psi}$ $=\int_{0}^{2\pi}e^{\bm{i}\theta}\int_{\infty}^{\infty}P(\theta,t,\omega)g(\omega)\;d\omega\;d\theta.$ (72)
@@ 3.1.3 The asymmetric Hopf model
for j=1,,Nj=1,...,N, where $z_{j}=x_{j}+\bm{i}y_{j}$ is the complex-valued signal for node $j$, $\omega_{j}$ is the intrinsic regional frequency, $a_{j}$ is a bifurcation parameter, $\bm{A}$ is the network of interactions and $\eta_{j}$ is additive noise that is uncorrelated in time and between nodes. In the absence of noise and interactions, for $a_{j}<0$, the dynamics converge to a stable fixed point. The system goes through a bifurcation at $a_{j}=0$ after which the dynamics then converge to the stable limit-cycle. The model can also be written in Cartesian coordinates,
@@
for j=1,,Nj=1,...,N, where, in both cases, $x_{j}(t)$ is considered to be the activity of region $j$ over time.